# Radiated energy fraction of SPI-induced disruptions at ASDEX Upgrade

### v1

@@ 4 Experimental values of $\text{W}_{\text{rad}}$ and $\text{f}_{\text{rad}}$
While the non-disruptive SPI injections have a radiated energy fraction below 30%, the fraction increases during the TQ, CQ, and VDE phase of disruptive injections to values of $\leq 40$% in total. As indicated by discharge #40657 in figure 7(a), the radiation of the full-sized, 8 mm diameter pellet in the pre-TQ phase is only slightly higher compared to the non-disruptive injections into H-mode plasmas. When adding the radiated energy of the TQ and also the VDE phase on top of the radiation at the fragment arrival, the difference in $\text{f}_{\text{rad}}$ between the non-disruptive ($<20$%) and disruptive ($\leq 40$%) illustrated in figure 7(b) is explained. For pure deuterium SPI induced disruptions, the radiation at fragment arrival and the disruption processes (TQ, VDE, etc.) are of comparable levels. The stochastisation of the field lines during the TQ might lead to the additional radiation as more heat from the core region is transported into the scrape-off layer (SOL) plasma, where a higher (background) impurity concentration is present, consequently increasing the radiation. Additionally, higher heat loads during the pre-TQ, TQ and CQ/VDE phase could lead to increased impurity release from the wall, increasing the radiated power. As a consequence, it is important to keep in mind, that while the radiated energy (fraction) is often used as a first indicator of the mitigation efficiency, it does not fully reflect the strain on the components induced by thermal loads, currents or forces. Despite 40% of the energy being radiated for pure deuterium SPI causing a disruption, thermal loads might still high.
@@ 4.1 Impact of the shatter head geometry on $\text{f}_{\text{rad}}$
The aim of this section is to compare the effect of the different shatter head geometries and fragment size distributions on the radiated energy fraction. The fragment size is heavily influenced by the normal velocity in the shatter head of angle $\alpha$: $v_{\perp}=v\cdot\sin(\alpha)$, where higher normal velocities typically lead to smaller mean fragment size [29, 30, 27]. To illustrate the effects, we start by deriving a heuristic fit function for $\text{f}_{\text{rad}}$ in the following. We use the simple definition of $\text{f}_{\text{rad}}$ as
$\text{f}_{\text{rad}}=\frac{\text{P}_{\text{rad}}}{\text{P}_{\text{rad}}+\text{P}_{\text{thFW}}}=\frac{1}{1+x},$ (10)
with the thermal energy flux to the wall $\text{P}_{\text{thFW}}$ and consequently $x=\text{P}_{\text{thFW}}/\text{P}_{\text{rad}}$. Physics of the plasma sheath at the plasma-wall interface reflect the following scaling for the heat fluxes onto the plasma facing components [57, 58]:
$\text{P}_{\text{thFW}}\propto\text{n}_{\text{e}}\cdot\text{T}_{\text{e}}^{3/2},$ (11)
@@
with a constant $B$ (containing all proportionality factors) or the radiation fit parameter $\texttt{a}=\text{B}\cdot\text{G}_{\text{rad}}(\text{T}_{\text{e}})$ as well as the assimilation fit parameter b. The parameter b expresses the theoretical maximum number of impurity atoms a given target plasma could assimilate in the case of optimal delivery.
@@
For 8 mm diameter pellets injected into the SPI H-mode plasmas, figure 8 shows the effect of the shatter head geometry with the fit function from equation (15). The fits were created with the scipy.optimize.curve_fit function and the error estimate via np.sqrt(np.diag(pcov)) as suggested in the scipy documentation [60]. In table 4 the fitting parameters from equation (15) (including errors) are given. Hereby, the assimilation parameter b was fixed to allow a better comparison between the shatter heads as the fit parameters are not orthogonal. The fixed factor $\texttt{b}=5.85\times 10^{19}$ was determined by the total fit  t for all three shatter head geometries (dashed black line in figure 8). Hereby, lower values of a will result in higher values of $\text{f}_{\text{rad}}$, hence are considered optimal.
@@
t (total; black) $1.67\times 10^{19}$ $0.13\times 10^{19}$ (7.8%) $5.85\times 10^{19}$ $0.66\times 10^{19}$ (11.3%)
1 ($25^{\circ}$ rect.) $2.01\times 10^{19}$ $0.16\times 10^{19}$ (8.0%) $5.85\times 10^{19\ddagger}$
2 ($25^{\circ}$ circ.) $1.78\times 10^{19}$ $0.15\times 10^{19}$ (8.4%) $5.85\times 10^{19\ddagger}$
3 ($12.5^{\circ}$ rect.) $1.53\times 10^{19}$ $0.05\times 10^{19}$ (3.3%) $5.85\times 10^{19\ddagger}$
@@
For high neon content (above $10^{21}$ neon atoms – $\geq$10neon), the shatter head geometry – and therefore the fragment size and velocity distribution – seems to only play a minor role (2–5% difference; within the uncertainties estimated in section 3.1) as shown in figure 8. It shows a larger impact (10–20% difference) on injections with small neon doping (around 0.085–1% neon for the 8 mm pellets i.e. $10^{19}$$10^{20}$ neon atoms) which were performed to investigate the effects of low neon doping on the plasmoid drift suppression. Hereby, the $12.5^{\circ}$ rectangular shatter head – producing the larger and faster fragments compared to its $25^{\circ}$ counter part – is at the top of the $\text{f}_{\text{rad}}$ distribution. The lowest radiated energy values are observed for the $25^{\circ}$ rectangular shatter head. The short, $25^{\circ}$ circular shatter head is mostly in between these two and the $\text{f}_{\text{rad}}$-value varies strongly, as its effective shatter angle is a function of pellet impact position inside the shatter head [27] shown in figure 2(d). This is also in line with the results for pure deuterium SPI presented by Jachmich et al [23, 61] for the material assimilation – which is crucial to increase the free electron density in the core to suppress runaway electron generation for future machines. Hereby, shallower shatter geometries seem beneficial for material assimilation. We observe the strongest impact of the SPI parameters (head geometry, pre-shattering pellet velocity, neon content, …) on the disruption behaviour (i.e. frad, CQ dynamics, …) for the low neon doping not only in the experiments [62] but also first DREAM simulations [63]. The cases with $\leq 10^{20}$ neon atoms in the pellet seem the most sensible to injection parameter changes. Probably, the reason for this is that for pure deuterium injection the radiation from the pellet materials is complemented by other, ill-controlled factors (such as wall conditions or impurities from the vessel), whereas above a critical number of neon atoms the neon line radiation dominates over the impact of different fragment delivery parameters.
@@
Figure 9 shows the radiated energy fraction as a function of the normal velocity component $\text{v}_{\perp}$ for the different neon concentrations used as colour-coding. Similar to the observed trend in figure 8, we can observe at low neon content – indicated as 0.17% in the figure – that lower values of $\text{v}_{\perp}$ (i.e. larger fragments) to lead to higher $\text{f}_{\text{rad}}$-values. This is beneficial for ITER, where the present SPI design choice foresees a $15.5^{\circ}$ angle for the shattering unit [10, 64].
@@
For high neon content (10% neon, $2\times 10^{21}$ neon atoms) – where $\text{f}_{\text{rad}}$ saturates – the difference in $\text{f}_{\text{rad}}$ is found within the uncertainty estimate, and thus a significant trend has not been observed. There are relatively few data points available for 10% neon pellets, hence further experiments will be necessary to verify a potential change in trend from low to high neon content.
@@ 5 Summary
Based on experimental results at ASDEX Upgrade [36] and JET [42] as well as JOREK simulations for AUG VDEs [8], the coupled energy fraction $\text{f}_{\text{c}}$ was estimated to 50%, necessary to calculate the radiated energy fraction $\text{f}_{\text{rad}}$. The total radiated energy $\text{W}_{\text{rad}}$ and radiated energy fraction $\text{f}_{\text{rad}}$ are a strong function of the neon content inside the pellet. While non-disruptive SPI shows radiated energy fractions of $\leq 20$%, the fraction increases to $\leq 40$% during the TQ and VDE phase of disruptive injections. Already with small neon doping of the pellets ($10^{19}$$10^{20}$ neon atoms, equivalent to 0.085%–1% neon inside 8 mm diameter pellets) – to study the ablation physics and plasmoid drift suppression – $\text{f}_{\text{rad}}$-values between 40–70% are observed. Increasing the neon content even further, the $\text{f}_{\text{rad}}$-curve saturates around 80%. Overall, SPI with the $12.5^{\circ}$, rectangular shatter head – producing large and fast fragments – caused the highest radiated energy fractions. The largest effect of the shatter head geometry was observed for low neon doping of $\approx$ 0.085–0.17% in the 8 mm diameter pellets ($\lesssim 2\times 10^{19}$ neon atoms), with an increase in radiated energy in the 10–20% range. Largest sensitivity to injection parameters was also found for this amount of neon doping in recent DREAM simulations [63]. The effect of the shatter head geometry on the radiated energy fraction for high neon concentrations (above $2\times 10^{21}$ neon atoms or 10% neon 8 mm pellets) was observed to be in the 2–5% range, which is within the present uncertainty estimate of inter-shot comparisons. Comparing injections of different normal velocity components (proxy for fragment size), large fragments seem beneficial for $\text{f}_{\text{rad}}$ in the case of neon doped pellets (below approximately $2\times 10^{20}$ neon atoms), while a clear trend can not be proven with the current uncertainties for higher neon concentrations. We observed that full-size, 8 mm diameter, pure deuterium pellets are sufficient to cause the H-mode plasmas to disrupt. In contrast, plasmas with injections of 4 mm pellets or shortened 8 mm pellets (with about half the nominal length) can typically recover after the injection. No dedicated, fine scan of the amount of injected deuterium to trigger the disruption had been performed, however, the threshold is roughly located around $1\times 10^{22}$ deuterium molecules in line with previous simulations results [50].

### v2

@@ 4 Experimental values of $\text{W}_{\text{rad}}$ and $\text{f}_{\text{rad}}$
While the non-disruptive SPI injections have a radiated energy fraction below 30%, the fraction increases during the TQ, CQ, and VDE phase of disruptive injections to values of $\leq 40$% in total. As indicated by discharge #40657 in figure 8(a), the radiation of the full-sized, 8 mm diameter pellet in the pre-TQ phase is only slightly higher compared to the non-disruptive injections into H-mode plasmas. When adding the radiated energy of the TQ and also the VDE phase on top of the radiation at the fragment arrival, the difference in $\text{f}_{\text{rad}}$ between the non-disruptive ($<20$%) and disruptive ($\leq 40$%) illustrated in figure 8(b) is explained. For 100% $\textrm{D}_{2}$ SPI induced disruptions, the radiation at fragment arrival and the disruption processes (TQ, VDE, etc.) are of comparable levels. The stochastisation of the field lines during the TQ might lead to the additional radiation as more heat from the core region is transported into the scrape-off layer (SOL) plasma, where a higher (background) impurity concentration is present, consequently increasing the radiation. Additionally, higher heat loads during the pre-TQ, TQ and CQ/VDE phase could lead to increased impurity release from the wall, increasing the radiated power. As a consequence, it is important to keep in mind, that while the radiated energy (fraction) is often used as a first indicator of the mitigation efficiency, it does not fully reflect the strain on the components induced by thermal loads, currents or forces. Despite 40% of the energy being radiated for 100% $\textrm{D}_{2}$ SPI causing a disruption, thermal loads might still be high.
@@ 4.1 Impact of the shatter head geometry on $\text{f}_{\text{rad}}$
The aim of this section is to compare the effect of the different shatter head geometries and fragment size distributions on the radiated energy fraction. The fragment size is heavily influenced by the normal velocity in the shatter head of angle $\alpha$: $v_{\perp}=v\cdot\sin(\alpha)$, where higher normal velocities typically lead to smaller mean fragment size [30, 31, 28]. To illustrate the effects, we start by deriving a heuristic fit function for $\text{f}_{\text{rad}}$ in the following. We use the simple definition
$\text{f}_{\text{rad}}=\frac{\text{P}_{\text{rad}}}{\text{P}_{\text{rad}}+\text{P}_{\text{con}}}=\frac{1}{1+x},$ (10)
with the thermal energy flux to the wall $\text{P}_{\text{con}}$ and consequently $x=\text{P}_{\text{con}}/\text{P}_{\text{rad}}$. Hereby, $\text{P}_{\text{con}}$ refers to the total power going to the wall from ions and electrons, accounting for their kinetic and thermal energy (convected & conducted; compare $\textrm{W}_{\textrm{con}}$ in figure 3). Physics of the plasma sheath at the plasma-wall interface reflect the following scaling for the heat fluxes onto the plasma facing components [60, 61]:
$\text{P}_{\text{con}}\propto\text{n}_{\text{e}}\cdot\text{T}_{\text{e}}^{3/2},$ (11)
@@
with a constant $B$ (containing all proportionality factors) or the radiation fit parameter $\texttt{a}=\text{B}\cdot\text{G}_{\text{rad}}(\text{T}_{\text{e}})$ as well as the assimilation fit parameter b. While the fit parameter a includes the temperature dependent radiation properties, parameter b expresses the theoretical maximum number of impurity atoms a given target plasma could assimilate in the case of optimal delivery. As the plasma has a finite energy, it cannot ablate an arbitrary amount of material from the fragments. Thus there must be a limit on the assimilated atoms within the plasma, which is resembled by b. Also there is a trade-off between TQ duration and the time it takes for the shards to go through the plasma, the shards can only ablate during this time with a given ablation rate. Note, the parameters a and b do not span an orthogonal basis, hence by keeping e.g. b as a fixed parameter, the open fit parameter a can be used to compare the fit more quantitatively. Therefore, the fit parameter b is not necessarily equal to the number of assimilated atoms directly.
@@
For 8 mm diameter pellets injected into the SPI H-mode plasmas, figure 9 shows the effect of the shatter head geometry with the fit function from equation (15). The fits were created with the scipy.optimize.curve_fit function and the error estimate via np.sqrt(np.diag(pcov)) as suggested in the scipy documentation [63]. In table 4 the fitting parameters from equation (15) (including errors) are given. Hereby, the assimilation parameter b was fixed to allow a better comparison between the shatter heads as the fit parameters are not orthogonal. The fixed factor $\texttt{b}=6.57\times 10^{19}$ was determined by the total fit  t for all three shatter head geometries (dashed black line in figure 9). Hereby, lower values of a will result in higher values of $\text{f}_{\text{rad}}$, hence are considered optimal.
@@
t (total; black) $1.81\times 10^{19}$ $1.23\times 10^{18}$ (6.8%) $6.57\times 10^{19}$ $6.70\times 10^{18}$ (10.2%)
1 ($25^{\circ}$ rect.) $2.17\times 10^{19}$ $1.54\times 10^{18}$ (7.1%) $6.57\times 10^{19\ddagger}$
2 ($25^{\circ}$ circ.) $1.91\times 10^{19}$ $1.50\times 10^{18}$ (7.9%) $6.57\times 10^{19\ddagger}$
3 ($12.5^{\circ}$ rect.) $1.67\times 10^{19}$ $4.79\times 10^{17}$ (2.9%) $6.57\times 10^{19\ddagger}$
@@
For high neon content (above $10^{21}$ neon atoms; $\textrm{f}_{\textrm{neon}}\geq 10\%$ neon), the shatter head geometry – and therefore the fragment size and velocity distribution – seems to only play a minor role (2–5% difference; within the uncertainties estimated in section 3.1) as shown in figure 9. It shows a larger impact (10–20% difference) on injections with small neon doping (around 0.085–1% neon for the 8 mm pellets i.e. $10^{19}$$10^{20}$ neon atoms) which were performed to investigate the effects of low neon doping on the plasmoid drift suppression. Hereby, the $12.5^{\circ}$ rectangular shatter head – producing the larger and faster fragments compared to its $25^{\circ}$ counter part – is at the top of the $\text{f}_{\text{rad}}$ distribution. The lowest radiated energy values are observed for the $25^{\circ}$ rectangular shatter head. The short, $25^{\circ}$ circular shatter head is mostly in between these two and the $\text{f}_{\text{rad}}$-value varies strongly, as its effective shatter angle is a function of pellet impact position inside the shatter head [28] shown in figure 2(d). This is also in line with the results for 100% $\textrm{D}_{2}$ SPI presented by Jachmich et al [23, 64] for the material assimilation – which is crucial to increase the free electron density in the core to suppress runaway electron generation for future machines. Hereby, shallower shatter geometries seem beneficial for material assimilation. We observe the strongest impact of the SPI parameters (head geometry, pre-shattering pellet velocity, neon content, …) on the disruption behaviour (i.e. frad, CQ dynamics, …) for the low neon doping not only in the experiments [40, 25] but also DREAM simulations [65, 66]. The cases with $\leq 10^{20}$ neon atoms in the pellet seem the most sensible to injection parameter changes. Probably, the reason for this is that for 100% $\textrm{D}_{2}$ injection the radiation from the pellet materials is complemented by other, ill-controlled factors (such as wall conditions or impurities from the vessel), whereas above a critical number of neon atoms the neon line radiation dominates over the impact of different fragment delivery parameters. Another potential explanation is the observed change of inside-out to outside-in cooling, which was observed for this parameter range in INDEX simulations by A. Patel [67]. For higher neon content, a cold front develops that lags slightly behind the pellets, causing an outside-in cooling of the plasma. Here, a variation in fragment size and velocity may cause different cooling/radiation properties.
@@
Figure 10 shows the radiated energy fraction as a function of the normal velocity component $\text{v}_{\perp}$ for the different neon concentrations used as colour-coding. Similar to the observed trend in figure 9, we can observe at low neon content – indicated as 0.17% in the figure – that lower values of $\text{v}_{\perp}$ (i.e. larger fragments) lead to higher $\text{f}_{\text{rad}}$-values. This is in line with the findings for material assimilation [23, 64], hence to maximize the assimilation of deuterium for ITER, the present SPI design choice foresees a $15.5^{\circ}$ angle for the shattering unit [10, 68].
Around 1.25% neon, frad seems to be independent of $\text{v}_{\perp}$, illustrated by the horizontal, yellow line in figure 9.
@@
For high neon content (10% neon, $2\times 10^{21}$ neon atoms) – where $\text{f}_{\text{rad}}$ saturates – the difference in $\text{f}_{\text{rad}}$ is found within the uncertainty estimate, and thus a significant trend has not been observed. A minor benefit of smaller fragments may be visible in the experiments [69, 70, 25] (see also figure 10) and AUG SPI JOREK simulations [72, 71]. There are relatively few data points available for the $25^{\circ}$ shatter heads with 10% neon pellets for a direct comparison to the $12.5^{\circ}$ head, thus further experiments will be necessary to verify a potential change in trend from low to high neon content.
At present, this apparent change in trend is not yet fully understood. However, we want to highlight two factors which may be contributing to this:
  • The potential change of inside-out towards outside-in cooling may take place in the low neon doping range [67], which could lead to different plasma radiation properties after the injections, strongly depending on the delivery of the material: large or small, slow or fast fragments and the degree of plasmoid drift suppression.
  • The fragment size is a non-linear function of $\text{v}_{\perp}$ as observed experimentally for the ASDEX Upgrade SPI system [28]. For high perpendicular velocities, the mean fragment size does not vary as much as the mean fragment velocity. This could cause a potential roll over effect, where the increased mean fragment velocity outweighs the reduced size.
In the case, the low neon doping would not be sufficient for a strong reduction in the plasmoid drift, the larger & faster fragments in the low neon range could lead to a better deposition of the material in the core region, hence increased material assimilation [23, 64] and radiation (see figures 9 and 10). For high neon content, the smaller fragments may enhance the outside-in cooling, however, further studies on the radiation effectiveness and $\text{f}_{\text{rad}}$ are currently prepared for the ASDEX Upgrade INDEX simulations.
@@ 5 Summary
Based on experimental results at ASDEX Upgrade [38] as well as JOREK simulations for AUG VDEs [8], the coupled energy fraction $\text{f}_{\text{c}}$ was estimated to 50%, necessary to calculate the radiated energy fraction $\text{f}_{\text{rad}}$. The total radiated energy $\text{W}_{\text{rad}}$ and radiated energy fraction $\text{f}_{\text{rad}}$ are a strong function of the neon content inside the pellet. While non-disruptive SPI shows radiated energy fractions of $\leq 20$%, the fraction increases to a total of $\leq 40$% after the TQ and VDE phase of disruptive injections. Already with small neon doping of the pellets ($10^{19}$$10^{20}$ neon atoms, equivalent to 0.085%–1% neon inside 8 mm diameter pellets) – to study the ablation physics and plasmoid drift suppression – $\text{f}_{\text{rad}}$-values between 40–70% are observed. Increasing the neon content even further, the $\text{f}_{\text{rad}}$-curve saturates around 80%. Overall, SPI with the $12.5^{\circ}$, rectangular shatter head – producing large and fast fragments – caused the highest radiated energy fractions. The largest effect of the shatter head geometry was observed for low neon doping of $\approx$ 0.085–0.17% in the 8 mm diameter pellets ($\lesssim 2\times 10^{19}$ neon atoms), with an increase in radiated energy in the 10–20% range. Largest sensitivity to injection parameters was also found for this amount of neon doping in recent DREAM simulations [65, 66]. The effect of the shatter head geometry on the radiated energy fraction for high neon concentrations (above $2\times 10^{21}$ neon atoms or 10% neon 8 mm pellets) was observed to be in the 2–5% range, which is within the present uncertainty estimate of inter-shot comparisons. Comparing injections of different normal velocity components (proxy for fragment size), large fragments seem beneficial for $\text{f}_{\text{rad}}$ in the case of neon doped pellets (below approximately $2\times 10^{20}$ neon atoms), while a clear trend can not be proven with the current uncertainties for higher neon concentrations. We observed that full-size, 8 mm diameter, 100% $\textrm{D}_{2}$ pellets are sufficient to cause the H-mode plasmas to disrupt. In contrast, plasmas with injections of 4 mm pellets or shortened 8 mm pellets (with about half the nominal length) can typically recover after the injection. No dedicated, fine scan of the amount of injected deuterium to trigger the disruption had been performed, however, the threshold is roughly located around $1\times 10^{22}$ deuterium molecules in line with previous simulations results [52].